OpenStudy (anonymous):

Quadratic Equation 1.)(2/x)^2-4=7 2.)1/2x+3/2x^2-5=4

OpenStudy (lgbasallote):

and the question is...?

OpenStudy (anonymous):

identify A,B,C,

OpenStudy (unklerhaukus):

\[\left(\frac{2}{x}\right)^2-4=7\]\[\frac{4}{x^2}=11\]\[4=11x^2\]\[11x^2-4=0\] now your first equation is of the form \[ax^2+by+c=0\] you can apply the quadratic formula \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]